# Finitistic dimensions over commutative DG-rings

### v3

@@ 1. Introduction
In this paper, we answer the analogous questions for commutative noetherian DG-rings. Throughout we work with non-positively graded commutative noetherian DG-rings of finite amplitude (in cohomological notation). We note that by [12, Theorem 0.2] and [26, Theorem 7.21], such DG-rings never have finite global dimension. Our first main result, which appears as 5.2 in the main body, provides an upper bound on the projective dimension of any bounded DG-module of finite flat dimension.
@@
Note that if $A$ is an ordinary ring (i.e., a DG-ring concentrated in one degree), then the above theorem encompasses the classic results of Bass [3] and Raynaud–Gruson [17]. However, this is where the similarity to the world of ordinary rings ends, for the inequalities cannot be improved upon. Infact, in 8.5 we show that is possible for any $d\geq 0$ and $n>0$, to construct commutative noetherian DG-rings $A$ and $B$ such that $\operatorname{amp}(A)=\operatorname{amp}(B)=n$ and $\dim(\mathrm{H}^{0}(A))=\dim(\mathrm{H}^{0}(B))=d$, but $\mathsf{FPD}(A)=d-n$ and $\mathsf{FPD}(B)=d$. Thus either bound can be achieved in the DG-setting, and moreover, this dichotomy occurs for any dimension and amplitude.
@@ Acknowledgements
The second-named author thanks Jan Trlifaj for helpful discussions. Isaac Bird, Liran Shaul and Jordan Williamson were supported by the grant GA ČR 20-02760Y from the Czech Science Foundation. Prashanth Sridhar was supported by the grant GA ČR 20-13778S from the Czech Science Foundation.
@@ 2.2. Homological dimensions
We now recall the central players of our study, which are homological dimensions of DG-modules over DG-rings. These definitions extend those made in [1] for unbounded complexes over discrete rings to the DG-setting.
@@
We note that an alternative proof of the projective case can be found at [15, Theorem 2.22]. As mentioned above, the homological dimensions behave very well with respect to reduction and coreduction, as the following corollary illustrates.
@@
The crucial fact about $\mathrm{Proj}(A)$, as illustrated in [15, Lemma 2.8], is that $\mathrm{H}^{0}(-):\mathrm{Proj}(A)\to\mathrm{Proj}(\mathrm{H}^{0}(A))$ is an equivalence of categories. In particular, $\mathrm{Proj}(A)$ is closed under coproducts, extensions, and summands. Combining these facts, the following lemma is trivial.
@@ 3.1. Bounds on suprema and infima
We first recall from [19, Proposition 3.3] and [26, Proof of Proposition 3.1] how reduction and coreduction interact with $\sup$ and $\inf$.
@@
This proves the first part of (i). For the second part, by applying 3.1 and [8, Lemma 2.1] we obtain isomorphisms
@@
By [25, tag 0310], it holds that $\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})\in\operatorname{Ass}(\mathrm{H}^{i}(M_{\bar{\mathfrak{p}}}))$, while the fact that $R_{\bar{\mathfrak{p}}}$ is a dualizing DG-module implies that $\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}})$ is a non-zero finitely generated $\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})$-module. Since $\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})\in\operatorname{Ass}(\mathrm{H}^{i}(M_{\bar{\mathfrak{p}}}))$, there is an injective map
$\operatorname{Hom}_{\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})}(\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}}),\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})/\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}}))\rightarrow\operatorname{Hom}_{\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})}(\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}}),\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}}))$
@@
We note the preceding result, for complexes over rings, can be found at [8, Proposition 2.2]. We obtain the following corollary.
@@ 3.2. Bounds for DG-modules of finite flat dimension
By 2.3 there is an equality $\operatorname{inj\,dim}_{A}(M)=\operatorname{inj\,dim}_{\mathrm{H}^{0}(A)}\left(\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),M)\right).$ Combining this with [1, Corollary 2.5.I], we have
@@
Over a commutative noetherian local ring $A$, it is particularly helpful to know the injective dimension of a dualizing complex $R$, should it exist. In this case, it is well known (see, for example [25, tag 0BUJ]) that $\operatorname{inj\,dim}_{A}(R)=\inf(R)+\text{dim}(A)$. The next result provides an analogy to this result in the DG-setting.
@@
The fact that $A$ has bounded cohomology implies by [26, Corollary 7.3] that $R$ is bounded. By 2.3 we know that
@@
By [26, Proposition 7.5], $\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),R)$ is a dualizing complex over $\mathrm{H}^{0}(A)$ and hence, combining the above with the aforementioned result for rings, we obtain that
@@
We may now give the main theorem of this section, which will be of significant use in what follows. This extends [8, Theorem 2.6] to the DG-setting.
@@ 4. Bounding projective dimension over local DG-rings
We construct an sppj resolution of $M$, in the sense of [15, Definition 2.17]. By 4.2(1) we may choose a map $f_{0}\colon P_{0}\to M$ in $\operatorname{\mathsf{D}}(A)$ such that $P_{0}\in\mathrm{Proj}(A)$ and $\mathrm{H}^{0}(f_{0})\colon\mathrm{H}^{0}(P_{0})\to\mathrm{H}^{0}(M)$ is surjective. As $P_{0}\in\operatorname{Proj}(A)$, it follows that $\sup(P_{0})=0$, and that $\operatorname{amp}(P_{0})\leq\operatorname{amp}(A)$ as $\mathrm{H}^{i}(P_{0})=\mathrm{H}^{i}(P_{0}\otimes^{\mathsf{L}}_{A}A)=\mathrm{H}^{0}(P_{0})\otimes_{\mathrm{H}^{0}(A)}\mathrm{H}^{i}(A)$ by Equation 2.6. We may extend the map $f_{0}$ to a distinguished triangle
@@
$\cong\operatorname{Ext}^{n-(k-\sum_{j=0}^{k-1}\sup(M_{-j}))}_{A}(M_{-(k-1)}[\sup(M_{-(k-1)}],K)$
$=\operatorname{Ext}^{n-a_{k-1}}_{A}(M_{-(k-1)}[\sup(M_{-(k-1)}],K).$
@@
Recall that for a faithfully flat map of commutative rings $A\to B$, we have $\operatorname{proj\,dim}_{B}(M\otimes^{\mathsf{L}}_{A}B)=\operatorname{proj\,dim}_{A}(M)$ for any $M\in\mathsf{D}(A)$ of finite projective dimension, see for example the remark following [5, Theorem 14.1.21]. Combining this observation with 2.3 and 4.8 one obtains the following lemma.
@@
shows that $\operatorname{flat\,dim}_{\widehat{A}}(M\otimes^{\mathsf{L}}_{A}\widehat{A})<\infty$. By [19, Proposition 7.21], the DG-ring $\widehat{A}$ has a dualizing DG-module, and hence $\operatorname{proj\,dim}_{\widehat{A}}(M\otimes^{\mathsf{L}}_{A}\widehat{A})\leq\dim(\mathrm{H}^{0}(\widehat{A}))-\inf(M\otimes^{\mathsf{L}}_{A}\widehat{A})$ by 4.6.
Since $A\to\widehat{A}$ is faithfully flat by [18, Corollary 4.6] and [20, Proposition 4.16], by applying 4.8 we have $\inf(M)=\inf(M\otimes^{\mathsf{L}}_{A}\widehat{A})$. Moreover, as $\mathrm{H}^{0}(\widehat{A})=\widehat{\mathrm{H}^{0}(A)}$ by [20, Proposition 4.16], we know that $\dim(\mathrm{H}^{0}(\widehat{A}))=\dim(\mathrm{H}^{0}(A))$. Since
@@ 5. Local-to-global for projective dimension
The aim of this section is to remove the local assumption from 4.1. There is an obstacle to performing a typical local-to-global argument, namely that, in general, projective dimension cannot be checked at localization at prime ideals. For example, there are flat modules $P$ over commutative noetherian rings $A$ of finite Krull dimension, such that $P$ is not projective over $A$, but $P_{\mathfrak{p}}$ is a projective $A_{\mathfrak{p}}$-module for all $\mathfrak{p}\in\operatorname{Spec}(A)$. Such examples exist even over rings as simple as $\mathbb{Z}$, see [6, Exercise 19.12] for an example.
@@
Let $F\cong N$ be a flat resolution, and let $Y=A^{\mathbb{W}}F$, the minimal semi-flat cotorsion replacement of $F$ over $A$, described in [16, Construction 3.3] and [16, Theorem 3.4]. Then by definition $Y\cong F\cong N$, and as explained in the proof of [16, Theorem 4.9], [16, (1.17)] gives the required filtration. ∎
We now turn to the first main theorem in this paper, which gives the global case of 4.1. This generalizes the results of Raynaud-Gruson [17, Theorem II.3.2.6] to the setting of DG-rings.
@@
There is nothing to prove if $\dim(\mathrm{H}^{0}(A))=\infty$, so assume that $\dim(\mathrm{H}^{0}(A))=d<\infty$. To show that $\operatorname{proj\,dim}_{A}(M)\leq\dim(\mathrm{H}^{0}(A))-\inf(M)$ we use 2.2. Thus, for any $N\in\operatorname{\mathsf{D}}^{0}(A)$ we must show that $\operatorname{Ext}^{i}_{A}(M,N)=0$ for all $i>\dim(\mathrm{H}^{0}(A))-\inf(M)$. By shifting if necessary, we can without loss of generality assume that $\sup(N)=\inf(N)=0$. As $N\cong\mathrm{H}^{0}(N)$, and as the latter is a $\mathrm{H}^{0}(A)$-module, we may consider $N$ as a $\mathrm{H}^{0}(A)$-module. We now proceed as in the proof of [16, Theorem 4.9], and use 5.1 to first replace $N$ by a flat resolution $F\to N$ in $\operatorname{\mathsf{D}}(\mathrm{H}^{0}(A))$, and then replace $F$ by $Y=A^{\mathbb{W}}F$, the minimal semi-flat cotorsion replacement of $N$ over $\mathrm{H}^{0}(A)$. Here, we have set
@@
By assumption, the leftmost term is zero, so in order for us to show that $\operatorname{Ext}^{i}_{A}(M,Y_{j-1})=0$, it is thus enough to show that $\operatorname{Ext}^{i}_{A}(M,Y_{j-1}/Y_{j})=0$. This was also concluded in [16, Theorem 4.9], and the above shows this remains true in this setting. As stated in 5.1, $Y_{j-1}/Y_{j}$ is a complex of $\mathrm{H}^{0}(A)$-modules which decomposes as a direct product $\prod_{\bar{\mathfrak{p}}\in W_{j-1}}T(\bar{\mathfrak{p}})$, where $T(\bar{\mathfrak{p}})$ is a complex of flat cotorsion $\mathrm{H}^{0}(A)$-modules with cosupport in $\{\bar{\mathfrak{p}}\}$. Because of this direct product decomposition, it is thus enough to show that $\operatorname{Ext}^{i}_{A}(M,T(\bar{\mathfrak{p}}))=0$. Since $T_{\bar{\mathfrak{p}}}$ is a complex of $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$-modules, we may consider it as an object of $\operatorname{\mathsf{D}}(A_{\bar{\mathfrak{p}}})$, which implies by adjunction that
@@
Since by 2.3 we know it is possible to reduce the computation of projective dimension over $A$ to a computation of projective dimension over $\mathrm{H}^{0}(A)$, the reader might wonder whether it is possible to use this reduction to deduce 5.2 directly from the corresponding result for complexes over rings. However, this is not possible in general. Even though the analogous result of 5.2 holds for complexes over rings (see for example [5, Corollary 8.5.18]), the only conclusion one can obtain is that if $M\in\operatorname{\mathsf{D}}^{\mathrm{b}}(A)$ has $\operatorname{flat\,dim}_{A}(M)<\infty$, then

### v4

@@ 1. Introduction
In this paper, we answer the analogous questions for commutative noetherian DG-rings. Throughout we work with non-positively graded commutative noetherian DG-rings of finite amplitude (in cohomological notation). We note that by [13, Theorem 0.2] and [27, Theorem 7.21], such DG-rings never have finite global dimension. Our first main result, which appears as 5.2 in the main body, provides an upper bound on the projective dimension of any bounded DG-module of finite flat dimension.
@@
Note that if $A$ is an ordinary ring (i.e., a DG-ring concentrated in one degree), then the above theorem encompasses the classic results of Bass [4] and Raynaud–Gruson [18]. However, this is where the similarity to the world of ordinary rings ends, for the inequalities cannot be improved upon. Infact, in 8.5 we show that it is possible for any $d\geq 0$ and $n>0$, to construct commutative noetherian DG-rings $A$ and $B$ such that $\operatorname{amp}(A)=\operatorname{amp}(B)=n$ and $\dim(\mathrm{H}^{0}(A))=\dim(\mathrm{H}^{0}(B))=d$, but $\mathsf{FPD}(A)=d-n$ and $\mathsf{FPD}(B)=d$. Thus either bound can be achieved in the DG-setting, and moreover, this dichotomy occurs for any dimension and amplitude.
@@ Acknowledgements
The second-named author thanks Jan Trlifaj for helpful discussions. The authors thank the referees for various suggestions that helped improving the paper. Isaac Bird, Liran Shaul and Jordan Williamson were supported by the grant GA ČR 20-02760Y from the Czech Science Foundation. Prashanth Sridhar was supported by the grant GA ČR 20-13778S from the Czech Science Foundation.
@@ 2.2. Homological dimensions
We now recall the central players of our study, which are homological dimensions of DG-modules over DG-rings. These definitions extend those made in [2] for unbounded complexes over discrete rings to the DG-setting. In the DG setting, they were studied in  [1].
@@
We note that an alternative proof of the projective case can be found at [16, Theorem 2.22]. As mentioned above, the homological dimensions behave very well with respect to reduction and coreduction, as the following corollary illustrates.
@@
The crucial fact about $\mathrm{Proj}(A)$, as illustrated in [16, Lemma 2.8], is that $\mathrm{H}^{0}(-):\mathrm{Proj}(A)\to\mathrm{Proj}(\mathrm{H}^{0}(A))$ is an equivalence of categories. In particular, $\mathrm{Proj}(A)$ is closed under coproducts, extensions, and summands. Combining these facts, the following lemma is trivial.
@@ 3.1. Bounds on suprema and infima
We first recall from [20, Proposition 3.3] and [27, Proof of Proposition 3.1] how reduction and coreduction interact with $\sup$ and $\inf$.
@@
This proves the first part of (i). For the second part, by applying 3.1 and [9, Lemma 2.1] we obtain isomorphisms
@@
By [26, tag 0310], it holds that $\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})\in\operatorname{Ass}(\mathrm{H}^{i}(M_{\bar{\mathfrak{p}}}))$, while the fact that $R_{\bar{\mathfrak{p}}}$ is a dualizing DG-module implies that $\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}})$ is a non-zero finitely generated $\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})$-module. Since $\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})\in\operatorname{Ass}(\mathrm{H}^{i}(M_{\bar{\mathfrak{p}}}))$, there is an injective map
$\operatorname{Hom}_{\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})}(\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}}),\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})/\bar{\mathfrak{p}}\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}}))\rightarrow\operatorname{Hom}_{\mathrm{H}^{0}(A_{\bar{\mathfrak{p}}})}(\mathrm{H}^{s}(R_{\bar{\mathfrak{p}}}),\mathrm{H}^{i}(M_{\bar{\mathfrak{p}}}))$
@@
We note the preceding result, for complexes over rings, can be found at [9, Proposition 2.2]. We obtain the following corollary.
@@ 3.2. Bounds for DG-modules of finite flat dimension
By 2.3 there is an equality $\operatorname{inj\,dim}_{A}(M)=\operatorname{inj\,dim}_{\mathrm{H}^{0}(A)}\left(\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),M)\right).$ Combining this with [2, Corollary 2.5.I], we have
@@
Over a commutative noetherian local ring $A$, it is particularly helpful to know the injective dimension of a dualizing complex $R$, should it exist. In this case, it is well known (see, for example [26, tag 0BUJ]) that $\operatorname{inj\,dim}_{A}(R)=\inf(R)+\text{dim}(A)$. The next result provides an analogy to this result in the DG-setting.
@@
The fact that $A$ has bounded cohomology implies by [27, Corollary 7.3] that $R$ is bounded. By 2.3 we know that
@@
By [27, Proposition 7.5], $\mathsf{R}\mathrm{Hom}_{A}(\mathrm{H}^{0}(A),R)$ is a dualizing complex over $\mathrm{H}^{0}(A)$ and hence, combining the above with the aforementioned result for rings, we obtain that
@@
We may now give the main theorem of this section, which will be of significant use in what follows. This extends [9, Theorem 2.6] to the DG-setting.
@@ 4. Bounding projective dimension over local DG-rings
We construct an sppj resolution of $M$, in the sense of [16, Definition 2.17]. By 4.2(1) we may choose a map $f_{0}\colon P_{0}\to M$ in $\operatorname{\mathsf{D}}(A)$ such that $P_{0}\in\mathrm{Proj}(A)$ and $\mathrm{H}^{0}(f_{0})\colon\mathrm{H}^{0}(P_{0})\to\mathrm{H}^{0}(M)$ is surjective. As $P_{0}\in\operatorname{Proj}(A)$, it follows that $\sup(P_{0})=0$, and that $\operatorname{amp}(P_{0})\leq\operatorname{amp}(A)$ as $\mathrm{H}^{i}(P_{0})=\mathrm{H}^{i}(P_{0}\otimes^{\mathsf{L}}_{A}A)=\mathrm{H}^{0}(P_{0})\otimes_{\mathrm{H}^{0}(A)}\mathrm{H}^{i}(A)$ by Equation 2.6. We may extend the map $f_{0}$ to a distinguished triangle
@@
$\cong\operatorname{Ext}^{n-(k-\sum_{j=0}^{k-1}\sup(M_{-j}))}_{A}(M_{-(k-1)}[\sup(M_{-(k-1)})],K)$
$=\operatorname{Ext}^{n-a_{k-1}}_{A}(M_{-(k-1)}[\sup(M_{-(k-1)})],K).$
@@
Recall that for a faithfully flat map of commutative rings $A\to B$, we have $\operatorname{proj\,dim}_{B}(M\otimes^{\mathsf{L}}_{A}B)=\operatorname{proj\,dim}_{A}(M)$ for any $M\in\mathsf{D}(A)$ of finite projective dimension, see for example the remark following [6, Theorem 14.1.21]. Combining this observation with 2.3 and 4.8 one obtains the following lemma.
@@
shows that $\operatorname{flat\,dim}_{\widehat{A}}(M\otimes^{\mathsf{L}}_{A}\widehat{A})<\infty$. By [20, Proposition 7.21], the DG-ring $\widehat{A}$ has a dualizing DG-module, and hence $\operatorname{proj\,dim}_{\widehat{A}}(M\otimes^{\mathsf{L}}_{A}\widehat{A})\leq\dim(\mathrm{H}^{0}(\widehat{A}))-\inf(M\otimes^{\mathsf{L}}_{A}\widehat{A})$ by 4.6.
Since $A\to\widehat{A}$ is faithfully flat by [19, Corollary 4.6] and [21, Proposition 4.16], by applying 4.8 we have $\inf(M)=\inf(M\otimes^{\mathsf{L}}_{A}\widehat{A})$. Moreover, as $\mathrm{H}^{0}(\widehat{A})=\widehat{\mathrm{H}^{0}(A)}$ by [21, Proposition 4.16], we know that $\dim(\mathrm{H}^{0}(\widehat{A}))=\dim(\mathrm{H}^{0}(A))$. Since
@@ 5. Local-to-global for projective dimension
The aim of this section is to remove the local assumption from 4.1. There is an obstacle to performing a typical local-to-global argument, namely that, in general, projective dimension cannot be checked at localization at prime ideals. For example, there are flat modules $P$ over commutative noetherian rings $A$ of finite Krull dimension, such that $P$ is not projective over $A$, but $P_{\mathfrak{p}}$ is a projective $A_{\mathfrak{p}}$-module for all $\mathfrak{p}\in\operatorname{Spec}(A)$. Such examples exist even over rings as simple as $\mathbb{Z}$, see [7, Exercise 19.12] for an example.
@@
Let $F\cong N$ be a flat resolution, and let $Y=A^{\mathbb{W}}F$, the minimal semi-flat cotorsion replacement of $F$ over $A$, described in [17, Construction 3.3] and [17, Theorem 3.4]. Then by definition $Y\cong F\cong N$, and as explained in the proof of [17, Theorem 4.9], [17, (1.17)] gives the required filtration. ∎
We now turn to the first main theorem in this paper, which gives the global case of 4.1. This generalizes the results of Raynaud-Gruson [18, Theorem II.3.2.6] to the setting of DG-rings.
@@
There is nothing to prove if $\dim(\mathrm{H}^{0}(A))=\infty$, so assume that $\dim(\mathrm{H}^{0}(A))=d<\infty$. To show that $\operatorname{proj\,dim}_{A}(M)\leq\dim(\mathrm{H}^{0}(A))-\inf(M)$ we use 2.2. Thus, for any $N\in\operatorname{\mathsf{D}}^{0}(A)$ we must show that $\operatorname{Ext}^{i}_{A}(M,N)=0$ for all $i>\dim(\mathrm{H}^{0}(A))-\inf(M)$. By shifting if necessary, we can without loss of generality assume that $\sup(N)=\inf(N)=0$. As $N\cong\mathrm{H}^{0}(N)$, and as the latter is a $\mathrm{H}^{0}(A)$-module, we may consider $N$ as a $\mathrm{H}^{0}(A)$-module. We now proceed as in the proof of [17, Theorem 4.9], and use 5.1 to first replace $N$ by a flat resolution $F\to N$ in $\operatorname{\mathsf{D}}(\mathrm{H}^{0}(A))$, and then replace $F$ by $Y=A^{\mathbb{W}}F$, the minimal semi-flat cotorsion replacement of $N$ over $\mathrm{H}^{0}(A)$. Here, we have set
@@
By assumption, the leftmost term is zero, so in order for us to show that $\operatorname{Ext}^{i}_{A}(M,Y_{j-1})=0$, it is thus enough to show that $\operatorname{Ext}^{i}_{A}(M,Y_{j-1}/Y_{j})=0$. This was also concluded in [17, Theorem 4.9], and the above shows this remains true in this setting. As stated in 5.1, $Y_{j-1}/Y_{j}$ is a complex of $\mathrm{H}^{0}(A)$-modules which decomposes as a direct product $\prod_{\bar{\mathfrak{p}}\in W_{j-1}}T(\bar{\mathfrak{p}})$, where $T(\bar{\mathfrak{p}})$ is a complex of flat cotorsion $\mathrm{H}^{0}(A)$-modules with cosupport in $\{\bar{\mathfrak{p}}\}$. Because of this direct product decomposition, it is thus enough to show that $\operatorname{Ext}^{i}_{A}(M,T(\bar{\mathfrak{p}}))=0$. Since $T_{\bar{\mathfrak{p}}}$ is a complex of $\mathrm{H}^{0}(A)_{\bar{\mathfrak{p}}}$-modules, we may consider it as an object of $\operatorname{\mathsf{D}}(A_{\bar{\mathfrak{p}}})$, which implies by adjunction that
@@
Since by 2.3 we know it is possible to reduce the computation of projective dimension over $A$ to a computation of projective dimension over $\mathrm{H}^{0}(A)$, the reader might wonder whether it is possible to use this reduction to deduce 5.2 directly from the corresponding result for complexes over rings. However, this is not possible in general. Even though the analogous result of 5.2 holds for complexes over rings (see for example [6, Corollary 8.5.18]), the only conclusion one can obtain is that if $M\in\operatorname{\mathsf{D}}^{\mathrm{b}}(A)$ has $\operatorname{flat\,dim}_{A}(M)<\infty$, then